Circular No. 6395

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

S/1995 S 5 B. Sicardy and F. Poulet, Paris Observatory; and J.-L. Beuzit and P.

Prado, European Southern Observatory, write: “Further to IAUC 6269, four images taken on 1995 Aug. 9 reveal the presence of an unresolved object. The observed separations, measured west of the planet’s center, are (typical accuracy 0′′.15): 1995 Aug. 9.24479 UT, 15′′.83; 9.24549, 15′′.99; 9.24618, 16′′.10; 9.24688, 16′′.14. We identified this object as S/1995 S 5, detected by the HST more than a day later (IAUC 6243; Nicholson et al. 1996, Science 272, 509). Combining the ESO data with the HST data, we derive the mean motion, 581.7 ± 2.1 deg/day; the longitude measured from the ascending node of Saturn’s equatorial plane on the earth’s equator (equinox J2000.0) at epoch 1995 Aug. 10.5 UT (at Saturn), 130^{o}.7; the estimated radius, 37 km; and the short-K magnitude at mean opposition, 16.0. The radius is estimated from the object’s integrated flux, assuming a spherical shape and the albedo of Saturn X (Janus). Assuming also a circular orbit, and including the effects of Saturn’s J_{2}, J_{4}, and J_{6}, the mean motion derived above yields a geometric orbital radius of 140 265 ± 330 km, closer to the F-ring radius (140 209 ± 4 km) than the previously derived value derived from the HST data alone (139 682 ± 370 km).”

GRO J1744–28 C. Kouveliotou, Universities Space Research Association (USRA); K.

Deal, P. Woods, M. Briggs, University of Alabama in Huntsville (UAH); B. A. Harmon, G. J. Fishman, Marshall Space Flight Center, NASA; J. van Paradijs, UAH and University of Amsterdam; M. H. Finger, USRA; and J. Kommers and W. H. G. Lewin, Massachusetts Institute of Technology, report: “BATSE is no longer detecting bursts or persistent emission from the bursting pulsar GRO J1744–28. A total of 3080 bursts were detected from this source (not corrected for earth occultation or other deadtime effects) between its onset on 1995 Dec. 2 until its cessation on 1996 May 3. During most of this 152-day interval, the burst rate was fairly constant at an average of ∼ 20 bursts/day, but after Apr. 27 the rate declined rapidly. The last detected event had a fluence of 1.2 × 10^{−7} erg cm^{−2}. The pulsing signal in the persistent emission was last detected on Apr. 26; after this date, the pulsed flux was below 7 × 10^{−10} erg cm^{−2} s^{−1} (r.m.s., between 20 and 40 keV). Between Apr. 30 and May 6, earth-occultation measurement yields a 3σ upper limit on the total flux above 20 keV of 2.7 × 10^{−9} erg cm^{−2} s^{−1}, or ∼ 150 mCrab.”

1996 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green